I'm lost, please help.
Graph the following functions. Include at least 3 (x,y) points for each one.
1. f(x) = 5x
2. f(x) = 4x+2
3. f(x) = (1/3)x
4. f(x) = log5x

Answer by RAY100(1637)

The most straightforward way to graph any function is to make a T box and plot the points.
After much practice ,,,,you can recoqnize patterns,,,but this does work
1) y=5x
X,,,,,,Y
______
0,,,,,0
1,,,,,5, ,,,,,,5*1=5
5,,,,25,,,,,,5*5=25
it is linear,,,(no powers over x^1),,,, the y intercept is (0),,,, and the slope is (+5)
remember,,,,y=mx+b,,,,with m=slope,,,,b=y intercept
2) y=4x+2
X,,,,,,,,,Y
0,,,,,,,,2
1,,,,,,,,6,,,,4(1)+2
-1/2,,,,0
it is linear,,,,y intercept is (+2),,,,,and the slope is (+4)
2)y=(1/3)x
X,,,,,Y
0,,,,0
3,,,,,1,,,,,,1/3 of 3
6,,,,,2,,,,,,1/3 of 6
it is linear,,,,y intercept is 0,,,,,slope is +1/3
4) y=log(5x)
X,,,,,,,Y
0,,,,,, (- infinity),,,,just use calc (5*0) log
1/5,,,0,,,,calculator (1/5*5=1)log
50,,,,2.4,,,calculator (5*50=250)log
log functions all start at minus infinity at x=0, rise to y=0 at log (1), and then rise slowly,in our case to y=2.4 at x=(50)